Carol has a wall in her bathroom that is 8 ft. by 15 ft. She has a mirror hanging on that wall that is 3 ft. by 5 ft. How many feet are not covered by the mirror?

Answer by richard1234(7193)

8*15 - 3*5 = 120 - 15 = 105 sq ft